# Optimal rates of approximation by shallow ReLUk neural networks and applications to nonparametric regression

### 21 Jun 2023 v2

@@ 2 Approximation rates for shallow neural networks
By the compactness of $\mathbb{B}^{d}$, $f_{n_{i}}$ convergence uniformly to $\widetilde{f}$. Hence $f=\widetilde{f}\in\mathcal{F}_{\sigma_{k}}(1)$ and $\mathcal{F}_{\sigma_{k}}(1)$ is closed in $L^{\infty}(\mathbb{B}^{d})$. ∎
The proof of Proposition 2.2 actually shows the approximation rate $\mathcal{O}(N^{-1/2})$ for the subclass $\mathcal{F}_{\sigma_{k}}(N,1)$. This rate can be improved if we take into account the smoothness of the activation function.
Lemma 2.3 ([Bach, 2017; Ma et al., 2022; Siegel and Xu, 2022]).
Let $2\leq p<\infty$, then there exists an $M\geq 1$ depending only on $k,p,d$ such that,
$\sup_{f\in\mathcal{F}_{\sigma_{k}}(1)}\inf_{f_{N}\in\mathcal{F}_{\sigma_{k}}(N,M)}\|f-f_{N}\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{pk+1}{pd}}.$
For $k\in\{0,1\}$ and $d>2k+1$, it holds that
$\sup_{f\in\mathcal{F}_{\sigma}(1)}\inf_{f_{N}\in\mathcal{F}_{\sigma}(N,M)}\|f-f_{N}\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{2k+1}{2d}}.$
Proof.
Using similar argument as Siegel and Xu [2022, Lemma 1], one can show that the parameterization map $\mathcal{P}_{\sigma_{k}}:\mathbb{S}^{d}\to L^{p}(\mathbb{B}^{d})$ given by $\mathcal{P}_{\sigma_{k}}(v)=\sigma_{k}((x^{\intercal},1)v)$ is of smoothness class $\alpha=k+1/p$ in the sense that, for each coordinate chart $(V,\phi)$ and $\xi\in L^{q}(\mathbb{B}^{d})$ where $1/p+1/q=1$, the function $f_{\xi}(u)=\langle\xi,\mathcal{P}_{\sigma_{k}}(\phi^{-1}(u))\rangle$ satisfies $\|f_{\xi}\|_{C^{r,\beta}(\phi(V))}\leq C\|\xi\|_{L^{q}(\mathbb{B}^{d})}$ for some constant $C<\infty$. By Siegel and Xu [2022, Theorem 2], if the parameterization map $\mathcal{P}_{\sigma_{k}}$ is of smoothness class $\alpha$, then there exists an $M\geq 1$ such that for any $f$ in the closure of $\cup_{N\in\mathbb{N}}\mathcal{F}_{\sigma_{k}}(N,1)$ in $L^{p}(\mathbb{B}^{d})$, we have
$\inf_{f_{N}\in\mathcal{F}_{\sigma_{k}}(N,M)}\|f-f_{N}\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{\alpha}{d}}.$
By Proposition 2.2, we can apply this bound to $f\in\mathcal{F}_{\sigma_{k}}(1)$.
The uniform bound for $k=0$ is obtained by [Ma et al., 2022, Theorem 4]. For $k=1$, the uniform bound is proved by Bach [2017, Proposition 1], who showed that approximating $f\in\mathcal{F}_{\sigma}(1)$ by finitely many neurons is essentially equivalent to the approximation of a zonoid by zonotopes [Matoušek, 1996]. The condition $d\geq 4$ is needed in order to use [Matoušek, 1996, Theorem 1]. We note that the proof also implies that one can take $M=3/2$. ∎
One can show that the rate $\mathcal{O}(N^{-\frac{1}{2}-\frac{2k+1}{2d}})$ in Lemma 2.3 is sharp for $p=2$ [Siegel and Xu, 2022]. Since the same rate holds in the supremum norm when $k=0,1$, we expect that it also holds for all $k\in\mathbb{N}_{0}$ and $2\leq p\leq\infty$. It seems that one needs to use techniques of geometric discrepancy theory to prove such kind of results [Bourgain et al., 1989; Matoušek, 1996].
@@
Combining Theorem 2.1 and Lemma 2.3, we can derive approximation rate of Hölder class $\mathcal{H}^{\alpha}$ for shallow neural network $\mathcal{F}_{\sigma_{k}}(N,M)$. Recall that we use the notation $a\lor b:=\max\{a,b\}$.
@@
Let $k\in\mathbb{N}_{0}$, $d\in\mathbb{N}$, $2\leq p<\infty$ and $\alpha>0$.
@@
    If $\alpha>(d+2k+1)/2$ or $\alpha=(d+2k+1)/2$ is an even integer, then there exists a constant $M$ depending on $k,p,d,\alpha$ such that
    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{pk+1}{pd}}.$
@@
    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{pk+1}{pd}}\sqrt{\log N}.$
@@
    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim N^{-\frac{\alpha}{d}\frac{d+2k+2/p}{d+2k+1}}\lor M^{-\frac{2\alpha}{d+2k+1-2\alpha}}.$
  4. (4)
    If $k\in\{0,1\}$ and $d>2k+1$, the bounds for $p=2$ also hold in the supremum norm. In particular, for $\alpha<(d+2k+1)/2$, we have
@@
    Thus, the rate $\mathcal{O}(N^{-\alpha/d})$ holds when $M\gtrsim N^{-(d+2k+1-2\alpha)/(2d)}$.
@@
We only present the proof for part (3), since other parts can be derived similarly. If $\alpha<(d+2k+1)/2$, then by Theorem 2.1, for any $h\in\mathcal{H}^{\alpha}$, there exists $g\in\mathcal{F}_{\sigma_{k}}(K)$ such that $\|h-g\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}$. Let $M_{0}$ be the constant in Lemma 2.3, then there exists $f\in\mathcal{F}_{\sigma_{k}}(N,KM_{0})$ such that $\|g-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim KN^{-\frac{1}{2}-\frac{pk+1}{pd}}$. If $M\geq N^{\frac{d+2k+1-2\alpha}{2d}\frac{d+2k+2/p}{d+2k+1}}$, we choose $K=N^{\frac{d+2k+1-2\alpha}{2d}\frac{d+2k+2/p}{d+2k+1}}/M_{0}$ then $f\in\mathcal{F}_{\sigma_{k}}(N,KM_{0})\subseteq\mathcal{F}_{\sigma_{k}}(N,M)$ and
$\|h-f\|_{L^{p}(\mathbb{B}^{d})}$ $\leq\|h-g\|_{L^{p}(\mathbb{B}^{d})}+\|g-f\|_{L^{p}(\mathbb{B}^{d})}$
$\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}+KN^{-\frac{d+2k+2/p}{2d}}$
$\lesssim N^{-\frac{\alpha}{d}\frac{d+2k+2/p}{d+2k+1}}.$
If $M\leq N^{\frac{d+2k+1-2\alpha}{2d}\frac{d+2k+2/p}{d+2k+1}}$, we choose $K=M/M_{0}$, then
$\|h-f\|_{L^{p}(\mathbb{B}^{d})}\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}+KN^{-\frac{d+2k+2/p}{2d}}\lesssim M^{-\frac{2\alpha}{d+2k+1-2\alpha}}.$
Combining the two bounds gives the desired result. By Lemma 2.3, if $k=0,1$ and $d>2k+1$, the bounds for $p=2$ also hold in the supremum norm. ∎
@@
We make some comments on the approximation rate for $\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$. As shown by Pinkus [1999, Corollary 6.10], the rate $\mathcal{O}(N^{-\alpha/d})$ in the $L^{2}$ norm is already known for $\alpha=1,2,\dots,(d+2k+1)/2$. For ReLU activation, the recent paper [Mao and Zhou, 2023] obtained the rate $\mathcal{O}(N^{-\frac{\alpha}{d}\frac{d+2}{d+4}})$ in the supremum norm. Corollary 2.4 shows that the rate $\mathcal{O}(N^{-\alpha/d})$ also holds in the supremum norm for ReLU and Heaviside activations. And more importantly, we also provide an explicit control on the network weights to ensure that such rate can be achieved, which is useful for estimating generalization errors (see Section 4). It is well-known that the optimal approximation rate for $\mathcal{H}^{\alpha}$ is $\mathcal{O}(N^{-\alpha/d})$, if we approximate $h\in\mathcal{H}^{\alpha}$ by a function class with $N$ parameters and the parameters are continuously depend on the target function $h$ [DeVore et al., 1989]. However, this result is not exactly applicable to neural networks, because we do not have guarantee that the parameters in the network depend continuously on the target function (in fact, this is not true for some constructions [Yarotsky, 2018; Lu et al., 2021; Yang et al., 2022]). Nevertheless, one can still prove that the rate $\mathcal{O}(N^{-\alpha/d})$ is optimal for shallow ReLUk neural networks by arguments based on pseudo-dimension as done in [Yarotsky, 2017; Lu et al., 2021; Yang et al., 2022].
@@
We describe the idea of proving approximation lower bound through pseudo-dimension by reviewing the result of Maiorov and Ratsaby [1999] (see also [Achour et al., 2022]). Recall that the pseudo-dimension ${\rm Pdim}\,(\mathcal{F})$ of a real-valued function class $\mathcal{F}$ defined on $\mathbb{B}^{d}$ is the largest integer $n$ for which there exist points $x_{1},\dots,x_{n}\in\mathbb{B}^{d}$ and constants $c_{1},\dots,c_{n}\in\mathbb{R}$ such that
@@
where $\mathcal{F}_{n}$ runs over all the classes in $L^{p}(\mathbb{B}^{d})$ with ${\rm Pdim}\,(\mathcal{F}_{n})\leq n$. They constructed a well-separated subclass of $\mathcal{H}^{\alpha}$ such that if a function class $\mathcal{F}$ can approximate this subclass with small error, then ${\rm Pdim}\,(\mathcal{F})$ should be large. In other words, the approximation error of any class $\mathcal{F}_{n}$ with ${\rm Pdim}\,(\mathcal{F}_{n})\leq n$ can be lower bounded. Consequently, they proved that
@@
By [Bartlett et al., 2019], we can upper bound the pseudo-dimension of shallow ReLUk neural networks as $n:=\,{\rm Pdim}\,(\mathcal{F}_{\sigma_{k}}(N,M))\lesssim N\log N$. Hence,
@@
which shows that the rate $\mathcal{O}(N^{-\alpha/d})$ in Corollary 2.4 is optimal in the $L^{2}$ norm (ignoring logarithmic factors). This also implies the optimality of Theorem 2.1 (otherwise, the proof of Corollary 2.4 would give a rate better than $\mathcal{O}(N^{-\alpha/d})$).
@@
Following the idea of Bach [2017], we first transfer the problem to approximation on spheres. Let us begin with a brief review of harmonic analysis on spheres [Dai and Xu, 2013]. For $n\in\mathbb{N}_{0}$, the spherical harmonic space $\mathbb{Y}_{n}$ of degree $n$ is the linear space that contains the restrictions of real harmonic homogeneous polynomials of degree $n$ on $\mathbb{R}^{d+1}$ to the sphere $\mathbb{S}^{d}$. The dimension of $\mathbb{Y}_{n}$ is $N(d,n):=\frac{2n+d-1}{n}\binom{n+d-2}{d-1}$ if $n\neq 0$ and $N(d,n):=1$ if $n=0$. Spherical harmonics are eigenfunctions of the Laplace-Beltrami operator:
@@
Let $\mathcal{P}_{n}:L^{2}(\mathbb{S}^{d})\to\mathbb{Y}_{n}$ denote the orthogonal projection operator. For any orthonormal basis $\{Y_{nj}:1\leq j\leq N(d,n)\}$ of $\mathbb{Y}_{n}$, the addition formula [Dai and Xu, 2013, Theorem 1.2.6] shows
@@
with normalization $P_{n}(1)=1$. Applying the Cauchy–Schwarz inequality to (3.1), we get $|P_{n}(t)|\leq 1$. For $n\neq 0$, $P_{n}(t)$ is odd (even) if $n$ is odd (even). Note that. for $d=1$ and $n\neq 0$, $N(d,n)=2$ and $P_{n}(t)$ is the Chebyshev polynomial such that $P_{n}(\cos\theta)=\cos(n\theta)$. We can write the projection $\mathcal{P}_{n}$ as
@@
Let $\varrho$ be the probability distribution with density $c_{d}(1-t^{2})^{(d-2)/2}$ on [1,1][-1,1], where the constant $c_{d}=(\int_{-1}^{1}(1-t^{2})^{(d-2)/2}dt)^{-1}=\omega_{d-1}/\omega_{d}$. For $f\in L^{1}(\mathbb{S}^{d})$ and $g\in L^{1}_{\varrho}([-1,1])$, define
@@
The convolution on the sphere satisfies Young’s inequality [Dai and Xu, 2013, Theorem 2.1.2]: for $p,q,r\geq 1$ with $p^{-1}=q^{-1}+r^{-1}-1$, it holds
@@
where the norm is the uniform norm when $r=\infty$. Observe that the projection $\mathcal{P}_{n}f=f*(N(d,n)P_{n})$ is a convolution operator with $\|N(d,n)P_{n}\|_{L^{\infty}([-1,1])}\leq N(d,n)$. Furthermore, for $g\in L^{1}_{\varrho}([-1,1])$, let $\widehat{g}(n)$ denote the Fourier coefficient of $g$ with respect to the Gegenbauer polynomials,
@@
By the Funk-Hecke formula, one can show that [Dai and Xu, 2013, Theorem 2.1.3]
@@
One of the key steps in our proof of Theorem 2.1 is the observation that functions of the form $f(u)=\int_{\mathbb{S}^{d}}\phi(v)\sigma_{k}(u^{\intercal}v)d\tau_{d}(v)$ are convolutions $\phi*\sigma_{k}$ with the activation function $\sigma_{k}\in L^{\infty}([-1,1])$. Bach [2017, Appendix D.2] has computed the Fourier coefficients $\widehat{\sigma_{k}}(n)$ explicitly. We summarize the result in the following.
@@
For even integers $\alpha=2s$, one can also use combinations of $T_{j\theta}$ and obtain [Rustamov, 1993; Ditzian, 2006]
@@
Another way to characterize the smoothness is through the $K$-functional. We first introduce the fractional Sobolev space induced by the Laplace-Beltrami operator. We say a function $f\in L^{p}(\mathbb{S}^{d})$ belong to the Sobolev space $\mathcal{W}^{\alpha,p}(\mathbb{S}^{d})$ if there exists a function in $L^{p}(\mathbb{S}^{d})$, which will be denoted by $(-\Delta)^{\alpha/2}f$, such that
@@
$K_{\alpha}(f,t)_{p}:=\inf_{g\in\mathcal{W}^{\alpha,p}(\mathbb{S}^{d})}\left\{\|f-g\|_{L^{p}(\mathbb{S}^{d})}+t^{\alpha}\|(-\Delta)^{\alpha/2}g\|_{L^{p}(\mathbb{S}^{d})}\right\}.$
It can be shown [Dai and Xu, 2013, Theorem 10.4.1] that the moduli of smoothness and the $K$-functional are equivalent:
@@
as the corresponding function class of $\mathcal{F}_{\sigma_{k}}(M)$ on $\mathbb{S}^{d}$. Abusing the notation, we will also denote $\gamma(g)=\inf_{\mu}\|\mu\|$ as the variation norm of $g\in\mathcal{G}_{\sigma_{k}}(M)$. The next proposition transfer our approximation problem on the unit ball $\mathbb{B}^{d}$ to approximation on the sphere $\mathbb{S}^{d}$.
Proposition 3.3.

### 09 Jan 2024 v3

@@ 2 Approximation rates for shallow neural networks
By the compactness of $\mathbb{B}^{d}$, $f_{n_{i}}$ converges uniformly to $\widetilde{f}$. Hence $f=\widetilde{f}\in\mathcal{F}_{\sigma_{k}}(1)$.
For $k=0$, we use the idea from [58, Lemma 3]. We can view $f_{n}$ as a Bochner integral $\int_{\mathbb{D}}i_{\mathbb{D}\to L^{2}(\mathbb{B}^{d})}d\mu_{n}$ of the inclusion map $i_{\mathbb{D}\to L^{2}(\mathbb{B}^{d})}$, where $\mathbb{D}:=\{g_{v}(x)=\sigma_{0}((x^{\intercal},1)v):v\in\mathbb{S}^{d}\}$. Notice that the set $\mathbb{D}\subseteq L^{2}(\mathbb{B}^{d})$ is compact, because the mapping $v\mapsto g_{v}$ is continuous. By Prokhorov’s theorem, there exists a weakly convergent subsequence $\mu_{n_{i}}\to\mu$. Let us denote $\widetilde{f}=\int_{\mathbb{D}}i_{\mathbb{D}\to L^{2}(\mathbb{B}^{d})}d\mu$, then $\widetilde{f}\in\mathcal{F}_{\sigma_{k}}(1)$ by viewing the Bochner integral as an integral over $\mathbb{S}^{d}$. If we choose a countable dense sequence $\{g_{j}\}_{j=1}^{\infty}$ of $L^{2}(\mathbb{B}^{d})$, then the weak convergence implies that
$\lim_{n_{i}\to\infty}\langle g_{j},f_{n_{i}}\rangle_{L^{2}(\mathbb{B}^{d})}=\left\langle g_{j},\widetilde{f}\right\rangle_{L^{2}(\mathbb{B}^{d})},$
for all $j$. The strong convergence $f_{n_{i}}\to f$ in $L^{\infty}(\mathbb{B}^{d})$ implies that the same equality for $f$ replacing $\widetilde{f}$. Therefore, $\langle g_{j},f\rangle_{L^{2}(\mathbb{B}^{d})}=\langle g_{j},\widetilde{f}\rangle_{L^{2}(\mathbb{B}^{d})}$ for all $j$, which shows $f=\widetilde{f}\in\mathcal{F}_{\sigma_{k}}(1)$. ∎
The proof of Proposition 2.2 actually shows the approximation rate $\mathcal{O}(N^{-1/2})$ for the subclass $\mathcal{F}_{\sigma_{k}}(N,1)$. This rate can be improved if we take into account the smoothness of the activation function. For ReLU activation, Bach [3, Proposition 1] showed that approximating $f\in\mathcal{F}_{\sigma}(1)$ by neural networks with finitely many neurons is essentially equivalent to the approximation of a zonoid by zonotopes [10, 37]. Using this equivalence, he obtained the rate $\mathcal{O}(N^{-\frac{1}{2}-\frac{3}{2d}})$ for ReLU neural networks. Similar idea was applied to the Heaviside activation in [32, Theorem 4], which proved the rate $\mathcal{O}(N^{-\frac{1}{2}-\frac{1}{2d}})$ for such an activation function. For ReLUk neural networks, the general approximation rate $\mathcal{O}(N^{-\frac{1}{2}-\frac{2k+1}{2d}})$ was established in $L^{2}$ norm by [57], which also showed that this rate is sharp. The recent work [55] further proved that this rate indeed holds in the uniform norm. We summarize their results in the following lemma.
Lemma 2.3 ([55]).
For $k\in\mathbb{N}_{0}$ and $d\in\mathbb{N}$, it holds that
$\sup_{f\in\mathcal{F}_{\sigma_{k}}(1)}\inf_{f_{N}\in\mathcal{F}_{\sigma_{k}}(N,1)}\|f-f_{N}\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{2k+1}{2d}}.$
@@
Combining Theorem 2.1 and Lemma 2.3, we can derive the rate of approximation by shallow neural network $\mathcal{F}_{\sigma_{k}}(N,M)$ for Hölder class $\mathcal{H}^{\alpha}$. Recall that we use the notation $a\lor b:=\max\{a,b\}$.
@@
Let $k\in\mathbb{N}_{0}$, $d\in\mathbb{N}$ and $\alpha>0$.
@@
    If $\alpha>(d+2k+1)/2$ or $\alpha=(d+2k+1)/2$ is an even integer, then there exists a constant $M$ depending on $k,d,\alpha$ such that
    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{2k+1}{2d}}.$
@@
    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma_{k}}(N,M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim N^{-\frac{1}{2}-\frac{2k+1}{2d}}\sqrt{\log N}.$
@@
    Thus, the rate $\mathcal{O}(N^{-\alpha/d})$ holds when $M\gtrsim N^{(d+2k+1-2\alpha)/(2d)}$.
@@
We only present the proof for part (3), since other parts can be derived similarly. If $\alpha<(d+2k+1)/2$, then by Theorem 2.1, for any $h\in\mathcal{H}^{\alpha}$, there exists $g\in\mathcal{F}_{\sigma_{k}}(K)$ such that $\|h-g\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}$. By Lemma 2.3, then there exists $f\in\mathcal{F}_{\sigma_{k}}(N,K)$ such that $\|g-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim KN^{-\frac{1}{2}-\frac{2k+1}{2d}}$. If $M\geq N^{\frac{d+2k+1-2\alpha}{2d}}$, we choose $K=N^{\frac{d+2k+1-2\alpha}{2d}}$ then $f\in\mathcal{F}_{\sigma_{k}}(N,K)\subseteq\mathcal{F}_{\sigma_{k}}(N,M)$ and
$\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}$ $\leq\|h-g\|_{L^{\infty}(\mathbb{B}^{d})}+\|g-f\|_{L^{\infty}(\mathbb{B}^{d})}$
$\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}+KN^{-\frac{d+2k+1}{2d}}\lesssim N^{-\frac{\alpha}{d}}.$
If $M\leq N^{\frac{d+2k+1-2\alpha}{2d}}$, we choose $K=M$, then
$\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim K^{-\frac{2\alpha}{d+2k+1-2\alpha}}+KN^{-\frac{d+2k+1}{2d}}\lesssim M^{-\frac{2\alpha}{d+2k+1-2\alpha}}.$
Combining the two bounds gives the desired result. ∎
@@
We make some comments on the approximation rate for $\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$. As shown by [48, Corollary 6.10], the rate $\mathcal{O}(N^{-\alpha/d})$ in the $L^{2}$ norm is already known for $\alpha=1,2,\dots,(d+2k+1)/2$. For ReLU activation, the recent paper [36] obtained the rate $\mathcal{O}(N^{-\frac{\alpha}{d}\frac{d+2}{d+4}})$ in the supremum norm. Corollary 2.4 shows that the rate $\mathcal{O}(N^{-\alpha/d})$ holds in the supremum norm for all ReLUk activations. And more importantly, we also provide an explicit control on the network weights to ensure that this rate can be achieved, which is useful for estimating generalization errors (see Section 4.2). It is well-known that the optimal approximation rate for $\mathcal{H}^{\alpha}$ is $\mathcal{O}(N^{-\alpha/d})$, if we approximate $h\in\mathcal{H}^{\alpha}$ by a function class with NN parameters and the parameters are continuously dependent on the target function $h$ [13]. However, this result is not directly applicable to neural networks, because we do not have guarantee that the parameters in the network depend continuously on the target function (in fact, this is not true for some constructions [65, 31, 63]). Nevertheless, one can still prove that the rate $\mathcal{O}(N^{-\alpha/d})$ is optimal for shallow ReLUk neural networks by arguments based on pseudo-dimension as done in [64, 31, 63].
@@
We describe the idea of proving approximation lower bounds through pseudo-dimension by reviewing the result of Maiorov and Ratsaby [33] (see also [1]). Recall that the pseudo-dimension ${\rm Pdim}\,(\mathcal{F})$ of a real-valued function class $\mathcal{F}$ defined on $\mathbb{B}^{d}$ is the largest integer $n$ for which there exist points $x_{1},\dots,x_{n}\in\mathbb{B}^{d}$ and constants $c_{1},\dots,c_{n}\in\mathbb{R}$ such that
@@
where $p\in[1,\infty]$ and $\mathcal{F}_{n}$ runs over all the classes in $L^{p}(\mathbb{B}^{d})$ with ${\rm Pdim}\,(\mathcal{F}_{n})\leq n$. They constructed a well-separated subclass of $\mathcal{H}^{\alpha}$ such that if a function class $\mathcal{F}$ can approximate this subclass with small error, then ${\rm Pdim}\,(\mathcal{F})$ should be large. In other words, the approximation error of any class $\mathcal{F}_{n}$ with ${\rm Pdim}\,(\mathcal{F}_{n})\leq n$ can be lower bounded. Consequently, they proved that
@@
By [6], we can upper bound the pseudo-dimension of shallow ReLUk neural networks as $n:=\,{\rm Pdim}\,(\mathcal{F}_{\sigma_{k}}(N,M))\lesssim N\log N$. Hence,
@@
which shows that the rate $\mathcal{O}(N^{-\alpha/d})$ in Corollary 2.4 is optimal in the $L^{p}$ norm (ignoring logarithmic factors). This also implies the optimality of Theorem 2.1 (otherwise, the proof of Corollary 2.4 would give a rate better than $\mathcal{O}(N^{-\alpha/d})$).
@@
Following the idea of [3], we first transfer the problem to approximation on spheres. Let us begin with a brief review of harmonic analysis on spheres [12]. For $n\in\mathbb{N}_{0}$, the spherical harmonic space $\mathbb{Y}_{n}$ of degree $n$ is the linear space that contains the restrictions of real harmonic homogeneous polynomials of degree $n$ on $\mathbb{R}^{d+1}$ to the sphere $\mathbb{S}^{d}$. The dimension of $\mathbb{Y}_{n}$ is $N(d,n):=\frac{2n+d-1}{n}\binom{n+d-2}{d-1}$ if $n\neq 0$ and $N(d,n):=1$ if $n=0$. Spherical harmonics are eigenfunctions of the Laplace-Beltrami operator:
@@
Let $\mathcal{P}_{n}:L^{2}(\mathbb{S}^{d})\to\mathbb{Y}_{n}$ denote the orthogonal projection operator. For any orthonormal basis $\{Y_{nj}:1\leq j\leq N(d,n)\}$ of $\mathbb{Y}_{n}$, the addition formula [12, Theorem 1.2.6] shows
@@
with normalization $P_{n}(1)=1$. Applying the Cauchy–Schwarz inequality to (3.1), we get $|P_{n}(t)|\leq 1$. For $n\neq 0$, $P_{n}(t)$ is odd (even) if $n$ is odd (even). Note that, for $d=1$ and $n\neq 0$, $N(d,n)=2$ and $P_{n}(t)$ is the Chebyshev polynomial such that $P_{n}(\cos\theta)=\cos(n\theta)$. We can write the projection $\mathcal{P}_{n}$ as
@@
Let $\varrho$ be the probability distribution with density $c_{d}(1-t^{2})^{(d-2)/2}$ on [1,1][-1,1], with the constant $c_{d}=(\int_{-1}^{1}(1-t^{2})^{(d-2)/2}dt)^{-1}=\omega_{d-1}/\omega_{d}$. For $f\in L^{1}(\mathbb{S}^{d})$ and $g\in L^{1}_{\varrho}([-1,1])$, define
@@
The convolution on the sphere satisfies Young’s inequality [12, Theorem 2.1.2]: for $p,q,r\geq 1$ with $p^{-1}=q^{-1}+r^{-1}-1$, it holds
@@
where the norm is the uniform one when $r=\infty$. Observe that the projection $\mathcal{P}_{n}f=f*(N(d,n)P_{n})$ is a convolution operator with $\|N(d,n)P_{n}\|_{L^{\infty}([-1,1])}\leq N(d,n)$. Furthermore, for $g\in L^{1}_{\varrho}([-1,1])$, let $\widehat{g}(n)$ denote the Fourier coefficient of $g$ with respect to the Gegenbauer polynomials,
@@
By the Funk-Hecke formula, one can show that [12, Theorem 2.1.3]
@@
One of the key steps in our proof of Theorem 2.1 is the observation that functions of the form $f(u)=\int_{\mathbb{S}^{d}}\phi(v)\sigma_{k}(u^{\intercal}v)d\tau_{d}(v)$ are convolutions $\phi*\sigma_{k}$ with the activation function $\sigma_{k}\in L^{\infty}([-1,1])$. [3, Appendix D.2] has computed the Fourier coefficients $\widehat{\sigma_{k}}(n)$ explicitly. We summarize the result in the following.
@@
For even integers $\alpha=2s$, one can also use combinations of $T_{j\theta}$ and obtain [50, 15]
@@
Another way to characterize the smoothness is through the $K$-functionals. We first introduce the fractional Sobolev space induced by the Laplace-Beltrami operator. We say a function $f\in L^{p}(\mathbb{S}^{d})$ belong to the Sobolev space $\mathcal{W}^{\alpha,p}(\mathbb{S}^{d})$ if there exists a function in $L^{p}(\mathbb{S}^{d})$, which will be denoted by $(-\Delta)^{\alpha/2}f$, such that
@@
$K_{\alpha}(f,t)_{p}:=\inf_{g\in\mathcal{W}^{\alpha,p}(\mathbb{S}^{d})}\left\{\|f-g\|_{L^{p}(\mathbb{S}^{d})}+t^{\alpha}\|(-\Delta)^{\alpha/2}g\|_{L^{p}(\mathbb{S}^{d})}\right\},\quad t>0.$
It can be shown [12, Theorem 10.4.1] that the moduli of smoothness and the $K$-functional are equivalent:
@@
as the corresponding function class of $\mathcal{F}_{\sigma_{k}}(M)$ on $\mathbb{S}^{d}$. Abusing the notation, we will also denote $\gamma(g)=\inf_{\mu}\|\mu\|$ as the variation norm of $g\in\mathcal{G}_{\sigma_{k}}(M)$. The next proposition transfers our approximation problem on the unit ball $\mathbb{B}^{d}$ to that on the sphere $\mathbb{S}^{d}$.
Proposition $3.$